kenzie126:

How are algebraic properties applied in the steps to solving equations and use them to solve ½x + 4 = 12?

JAIDENGYMNASTIC:

FIRST OF ALLL ..... WELCOMEE:)

JAIDENGYMNASTIC:

1/2*x+4=12 | -4 1/2*x=12-4 x/2=8 | *2 2x/2=8*2 x=16

kenzie126:

do i have to write down the steps

JAIDENGYMNASTIC:

you would want to so you can understand so yes